Find the equation of the line tangent to the given circle and passing through the given point
1. x^2+y^2= 26 (4,6)

Answer by Alan3354(69443)

Find the equation of the line tangent to the given circle and passing through the given point
1. x^2+y^2= 26 (4,6)
---------------------------
There are 2 tangent lines thru the point.
The point, the tangent points and the center make right triangles.
---
Find the distance from the point to the center of the circle
d = sqrt(52) = the hypotenuse
The radius is one side of the triangle, length = sqrt(26)
Use Pythagoras to find the distance from the point to the
tangent points, = sqrt(26)
--------------
The tangent points are the intersections of the given circle and a circle of radius sqrt(26) centered at (4,6)
-------------
Find the 2 intersections.
----
The 2 lines are the lines thru the tangent point and (4,6)
=================
email via the Thank You note for help, or to check your work.